### v1

@@ 1 Introduction
Our main theoretical results highlight a novel incentive to exacerbate societal divisions. In the conclusive sections of the paper we discuss examples of how this might occur, relating the main themes of the previous sections to the growing literature on zero-sum thinking (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023), common interest vs conflict of interest policy issues (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023), and identity politics (e.g. Shayo,, 2009; Bonomi et al.,, 2021; Gennaioli and Tabellini,, 2023). First, our discussion suggests that politicians might benefit if voters overestimate the zero-sum component of politics and perceive conflict of interest political issues as more salient than common interest policies, as this causes disagreement between opposite-minded voters to intensify.
@@
Our last application concerns party identity, modeled building on (Bonomi et al.,, 2021). We show that prompting voters to identify with the party they are more likely to vote for is not only valuable but also “efficient” for politicians, in that it distorts all voters’ bliss points in the direction of the fastest payoff improvement for candidates. These results complement those of Gennaioli and Tabellini, (2023), who show how competing parties might leverage other types of social identities for their electoral advantage. We also illustrate that politicians’ differentiation incentives can interact with voters’ party identity, with a feedback loop between party divergence and voter conflict that can generate a growth in polarization over time. Finally, we discuss how changing some features of the electoral rule might curb party divergence and the incentives to divide voters.
@@ 2 Model
The candidates of each party simultaneously announce their platforms $x_{A},x_{B}\in\mathbb{R}^{K}$. After the platforms are announced, but before citizens cast their vote, a popularity shock in favor of the candidate of party $B$, $\epsilon\in\mathbb{R}$, is realized. The shock has a continuous cumulative distribution $F$, and captures random events that influence voters’ relative assessment of party $B$, including, for instance, new information about the state of the economy, political scandals, or the effect of a candidate’s personality traits and characteristics on their appeal to voters. We assume that voters vote sincerely, taking into account $x_{A}$, $x_{B}$, and the popularity shock $\epsilon$. In particular, for each $x_{A},x_{B}$, let $\Delta_{i}(x_{A},x_{B})=v(x_{A},x_{i})-v(x_{B},x_{i})$. Each voter of type $i$ will vote for party $A$ if and only if $\Delta_{i}(x_{A},x_{B})\geq\epsilon$, which holds with probability $\pi_{i}(x_{A},x_{B})=F(\Delta_{i}(x_{A},x_{B}))$. As standard in probabilistic voting models, we let the support of $F$ be wide, so that cases where $\pi_{i}\in\{0,1\}$ can be ignored in the equilibrium analysis. [^7] Formally, for $x_{\lor}=(\max_{i\in I}{x^{k}_{i}})_{k\leq K}$ and $x_{\land}=(\min_{i\in I}{x^{k}_{i}})_{k\leq K}$, we assume that both $v(x_{\land},x_{\lor})-v(x_{\lor},x_{\lor})$ and $v(x_{\lor},x_{\lor})-v(x_{\land},x_{\lor})$ are in the support of $F$.
@@
Part (i) requires that rents increase in the number of votes. [^8] Proposition 2.1 would generalize to the case where $R$ is weakly increasing on $[0,\frac{1}{2}]$ and not constant on $[0,\frac{1}{2})$. In that case, party divergence would obtain for some distributions of voter types. We do not assume continuity: this allows, for instance, for jumps in the office rents as a party’s vote share meets institutional thresholds that grant agenda-setting power, the power to block legislation or entitlement to electoral reimbursements. In particular, the general setup allows for a party to gain a sizable share of power (and political rents) once it obtains the absolute majority of votes. An alternative interpretation of this setup (proposed, e.g., by Glaeser et al.,, 2005), is that voters are making a campaign contribution decision rather than a voting decision, so that parties’ rents naturally grow with their supporters. [^9] Under this interpretation, the size $s_{i}$ of voter group $i$ would already account for the heterogeneity in voters’ budget devoted to political contributions. Part (ii) imposes the restriction that aggregate rents are constant. Our second main assumption is that parties’ utility from rents is increasing and concave. Formally, let the payoff function of each party be $U:\mathbb{R}\to\mathbb{R}$. The following holds.
@@
Let $\nu:[0,1]\to\mathbb{R}$, $\nu=U\circ R$, be the function mapping a party’s vote share to its rents, and without any loss of generality, let payoffs be normalized so that $\nu(1)-\nu(0)=1$. The following holds when assumptions 1 and 2 are combined.
@@ 2.1 Party Divergence
In every pure strategy equilibrium, candidates offer different policy platforms.
@@ 3 Unidimensional Conflict
We start by analyzing the case where $K=1$, so that parties compete on a single policy dimension. [^12] In this case, dimension indexes $k$ are redundant, and omitted. Let $\epsilon\sim U(-\phi,\phi$), so that the shock is uniformly distributed, and assume that $h^{k}(|x-x_{i}|)=-\alpha(x^{k}-x^{k}_{i})^{2}$, where $\alpha>0$ captures how important policies are for individual payoffs and, without loss of generality, we set $\alpha=1$. With one policy dimension, we can think of voter heterogeneous preferences as capturing the classical economy-based conflict of interest between right-wing and left-wing voters, as in the following simple example.
@@
First, we show that the game admits a pure strategy equilibrium, and there are exactly two platforms that can be offered in such an equilibrium. One policy lies to the left of the risk-neutral equilibrium policy $x^{RN}$. The other lies to the right of such policy. For the next propositions, assume that voters are ordered so that $x_{1}\leq x_{2}\leq...\leq x_{N}$, and let $x_{m}$ be the median voter type. If the median voter does not exist, define $x_{m}$ to be the average of the two “most median” voter types.
Proposition 3.1 (Unidimensional Equilibrium)
The game admits a pure strategy equilibrium. Moreover, there exist two platforms, $\underaccent{\bar}{x},\bar{x}\in\mathbb{R}$ with $x_{1}<\underaccent{\bar}{x}<\bar{x}<x_{N}$, such that, in every pure strategy equilibrium, one candidate offers $\underaccent{\bar}{x}$ while the other candidate offers $\bar{x}$. It has
@@
where $w_{i},\omega_{i}\in(0,1)$, $\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=\sum_{i=1}^{N}\bar{w}_{i}=1$, and $\bar{w}_{i}>(<)\text{ }\underaccent{\bar}{w}^{i}$ if $x_{i}>(<)\text{ }x_{m}$, where $x_{m}$ is the median voter type. For each $i$, the equilibrium weights $\bar{w}_{i}$ and $\underaccent{\bar}{w}_{i}$ that parties attribute to voter group $i$ are proportional to their payoff increase from winning the votes of such group. Finally, the equilibrium expected payoff of each party is proportional to the (squared) distance between the two equilibrium platforms $(\bar{x}-\underaccent{\bar}{x})^{2}$, up to a constant term.
@@ 3.1 Voter Alienation and Polarization
In other words, a right-wing candidate benefits from attracting voter type $i$ if a change in $i$’s views on redistribution pushes her to support the political right and also increases the equilibrium payoff of the right-wing party. Conversely, the right-wing party benefits from alienating voter type $i$ if an increase in $i$’s support for redistribution makes the right-wing party better off in equilibrium, despite increasing left-wing support from group $i$.
At this point, one could be skeptical that a candidate ever benefits from alienating a group of voters, even if it means that the rest of the voter distribution remains unchanged and that group of voters will be less likely to support the party in the new equilibrium. Surprisingly, our next result suggests that each party always has the incentive to alienate at least one voter group: the most extremist supporters of the rival candidate.
@@ 4 Multidimensional Conflict
Let us now consider $K>1$ policies, so that platforms $x_{A}$ and $x_{B}$ are vectors of $\mathbb{R}^{K}$. While we still denote a voter distribution by $(\mathbf{x},\mathbf{s})$, note that now a type $x_{i}$ is also a vector of $\mathbb{R}^{K}$. The analysis is carried out under the same assumptions of section 3, with some additional restrictions. First, we require $\nu$ to be strictly concave. Strict concavity guarantees that party best responses never fall at points where payoffs are non-differentiable, which greatly simplifies the analysis. [^17] The assumption can be relaxed and, in particular, we can allow for a jump in payoff when parties reach the majority of votes. Second, for some of the main results of the section we restrict the attention to symmetric voter distribution, which we prove to admit a pure strategy Nash equilibrium. [^18] In the more general formulation of the game, the existence of a mixed strategy equilibrium is guaranteed by the Glicksberg theorem, but the existence of a pure strategy equilibrium is not, because the utility function of a party is not quasi-concave in that party’s strategy.
@@ 4.1 Equilibrium Analysis
The equilibrium analysis presents some technical challenges. First, parties’ expected payoffs are not differentiable at platform profiles $(x_{A},x_{B})$ such that some voter group $i$ is indifferent between voting for party $A$ or party $B$. Second, a party’s expected payoff is not quasi-concave in that party platform – best responses are not continuous in the competitors’ platforms. Third, the game is not finite, in the sense that the strategy set is non-finite. These features complicate the search for pure strategy equilibria, as even their existence is not guaranteed by standard results. The first step of the equilibrium analysis is introducing the notion of local equilibrium of the election game.
@@ 4.1.1 Algorithmic Solutions
Notes. Panel (a) plots a non-symmetric bidimensional voter distribution and the associated local equilibria where party payoffs. Panel (b) plots a symmetric bidimensional voter distribution and the associated local equilibria. In both panels: (i) ”x” markers represent voter types’ ideal policies, and each voter group is equally sized; (ii) two circles of the same color represent a pair of local equilibrium platforms; (iii) we assume that party’s payoff as a function of vote shares is $\nu(s)=\ln(\frac{1}{2}+s)$.
Figure 1 panel (a) reports a voter distribution over a bidimensional policy space, as well as the set of local equilibria, found applying the algorithm outlined in the previous paragraph. Lemma 4.1 tells us that among the existing local equilibria, the one identified by the two green dots is associated to the highest party expected payoff since the platforms proposed by the two politicians differ the most. In panel (b) we report the voter types and local equilibria for a symmetric voter distribution. In this case, the local equilibrium in blue is also the Nash equilibrium where parties expect to benefit the most, as per the results of proposition 4.1.
@@ 4.2 Divide and Diverge
Figure 2 reports the direction in which voter opinions should change for parties’ equilibrium payoffs to increase the fastest, for the two local equilibria of figure 1, panel (a). The red arrows represent this direction for the best local equilibrium (identified by the green circles), while the blue arrows identify the direction for the other local equilibrium (identified by the white circles). In both cases, the change makes supporters of party $A$ on average more different from supporters of party $B$.
@@ 5.1 Common Interest vs Conflict of Interest Issues
A recent stream of literature in political science and economics has stressed how politicians tend to focus on divisive, zero-sum, issues over common interest issues, even if the latter entail policies that could benefit a larger share of voters (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023). This phenomenon is potentially amplified by the fact that, on the one hand, voters seem to be prone to social and political “zero-sum thinking,” the perception that gains from one social group must necessarily come at the expense of another (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023), and often lack the information to assess the “common value” nature of policies (Matsusaka and Kendall,, 2022), while politicians’ rhetoric often paints politics in a zero-sum fashion. Our simple model can provide a new perspective on why politicians might have an incentive to focus on more divisive issues, stress the zero-sum component of policy-making, and keep voters informed on zero-sum issues and uninformed about common interest policies.
@@ 6 Some Words on Welfare
so that, when $R\to\bar{R}$, $\bar{x}\to x^{o}$ and $\underaccent{\bar}{x}\to x^{o}$. This result is captured by the following remark.
@@ A.1 Proofs
By proposition 2.1, we know that for any platform $x\in\mathbb{R}$, $x$ is never the best response to $x$. Fix a distribution $(\mathbf{x},\mathbf{s})$ and order $i,...,N$ so that $x_{1}\leq x_{2}\leq...\leq x_{N}$.
Fix $p\in\{A,B\}$ and let $x_{-p}\in\mathbb{R}$. As a first step we show that, if $x_{p}$ is a best response to $x_{-p}$, then either $x_{p}>x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$, for $\bar{w}_{i}=\frac{\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})}{\nu(1)-\nu(0)}$, or $x_{p}<x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}$, for $\underaccent{\bar}{w}_{i}=\frac{\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})}{\nu(1)-\nu(0)}$.
Assume that $x_{p}$ is a best response to $x_{-p}$. Given that $x_{p}\neq x_{-p}$ there are two cases. The first is $x_{p}>x_{-p}$. Subject to the constraint $x_{p}>x$, and fixing $x_{-p}$,$V_{p}$ is strictly concave in $p$’s platform $x$. Hence, $x_{p}$ it must satisfy the first-order condition $\frac{\partial V_{p}}{\partial x_{p}}=0$. When $p=A$, the first order condition is
@@
In order to show that a pure strategy Nash equilibrium exists, we need to show that $\bar{x}$ and $\underaccent{\bar}{x}$ are the best responses to each other. It suffices to show that $\underaccent{\bar}{x}<\bar{x}$, as we argue below. First, by proposition 2.1, $\underaccent{\bar}{x}$ is not the best response to $\underaccent{\bar}{x}$ and $\bar{x}$ is not the best response to $\bar{x}$. Hence, by the previous step of the proof, $\underaccent{\bar}{x}$ is the only possible best response to $\bar{x}$ and $\bar{x}$ is the only possible best response to $\underaccent{\bar}{x}$. Second, note that, for each $p=A,B$, $V_{p}$ is continuous, increasing in $x_{p}$ when $x_{p}<\underaccent{\bar}{x}$, decreasing in $x_{p}$ when $x_{p}>\bar{x}$. Hence, given the strict concavity of $V_{p}$ in $x_{p}$ when $x_{p}\neq x_{-p}$, $x_{p}$ is the best response to $\underaccent{\bar}{x}$ if $x_{p}>\underaccent{\bar}{x}$ and it satisfies the FOC of the respective maximization problem. The FOC requires $x_{p}=\bar{x}$. Similarly, $x_{p}$ is the best response to $\bar{x}$ if $x_{p}<\bar{x}$ and it satisfies the FOC of the respective maximization problem. The FOC requires $x_{p}=\underaccent{\bar}{x}$. Hence, $\underaccent{\bar}{x}<\bar{x}$ is a sufficient condition for $\bar{x}$ and $\underaccent{\bar}{x}$ to be the best responses to each other.
We now show $\underaccent{\bar}{x}<\bar{x}$. First, note that since $\underaccent{\bar}{w}_{i},\bar{w}_{i}\in[0,1]$ for $i=1,...,N$, and $\sum_{i=1}^{N}\bar{w}_{i}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=1$, $(\bar{w}_{i},x_{i})_{i\in I}$ and $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$ are probability distributions.
Second, $(\bar{w}_{i},x_{i})_{i\in I}$ strictly first order dominates $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$, as we now show. Since the probability distributions share support $\{x_{1},...,x_{N}\}$, it is sufficient to show that, for each $i=1,...,N$
@@
Having proved strict first order stochastic dominance, note that $\bar{x}$ is the expected value of $(\bar{w}_{i},x_{i})_{i\in I}$ and $\underaccent{\bar}{x}$ is the expected value of $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$. Strict first order stochastic dominance implies $\bar{x}>\underaccent{\bar}{x}$. This concludes the proof that the pure strategy equilibrium exists, and that only $\bar{x}$ and $\underaccent{\bar}{x}$ will be played in such equilibrium.
Next, note that for each $i\in I$, it has $\bar{w}_{i}=\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})$ and $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$. The former weight represents the payoff gain of the party offering $\bar{x}$ when she wins type $i$. Voter group $i$ will vote for the party offering $\bar{x}$ only if all $j>i$ – that is, a share $\bar{s}_{i+1}$ of the distribution, also vote for that party. If voter group $i$ also votes for the party offering $\bar{x}$, the vote share of that party increases from $\bar{s}_{i+1}$ to $\bar{s}_{i}=\bar{s}_{i+1}+s_{i}$, with an increase in payoff equal to $\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})=\bar{w}_{i}$. Using the same argument, it is easy to show that $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$ is the payoff gain of the party offering $\underaccent{\bar}{x}$ from winning voter group $i$. Note that, by remark 1, it has

### v2

@@ 1 Introduction
Our main theoretical results highlight a novel incentive to exacerbate societal divisions. In the conclusive sections of the paper we discuss examples of how this might occur, relating the main themes of the previous sections to the growing literature on zero-sum thinking (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023; Ali et al.,, 2024), common interest vs conflict of interest policy issues (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023), and identity politics (e.g. Shayo,, 2009; Bonomi et al.,, 2021; Gennaioli and Tabellini,, 2023). First, our discussion suggests that politicians might benefit if voters overestimate the zero-sum component of politics and perceive conflict of interest political issues as more salient than common interest policies, as this causes disagreement between opposite-minded voters to intensify.
@@
Our last application concerns party identity, modeled building on Bonomi et al., (2021). We show that prompting voters to identify with the party they are more likely to vote for is not only valuable but also “efficient” for politicians, in that it distorts all voters’ bliss points in the direction of the fastest payoff improvement for candidates. These results complement those of Gennaioli and Tabellini, (2023), who show how competing parties might leverage other types of social identities for their electoral advantage. We also illustrate that politicians’ differentiation incentives can interact with voters’ party identity, with a feedback loop between party divergence and voter conflict that can generate a growth in polarization over time. Finally, we discuss how changing some features of the electoral rule might curb party divergence and the incentives to divide voters.
@@ 2 Model
The candidates of each party simultaneously announce their platforms $x_{A},x_{B}\in\mathbb{R}^{K}$. After the platforms are announced, but before citizens cast their vote, a popularity shock in favor of the candidate of party $B$, $\epsilon\in\mathbb{R}$, is realized. The shock has a continuous cumulative distribution $F$, and captures random events that influence voters’ relative assessment of party $B$, including, for instance, new information about the state of the economy, political scandals, or the effect of a candidate’s personality traits and characteristics on their appeal to voters. We assume that voters vote sincerely, taking into account $x_{A}$, $x_{B}$, and the popularity shock $\epsilon$ [^7] Note that sincere voting is without loss of generality with atomistic voters, a scenario allowed in our setup. In fact, we only assume a finite number of voter types. In particular, for each $x_{A},x_{B}$, let $\Delta_{i}(x_{A},x_{B})=v(x_{A},x_{i})-v(x_{B},x_{i})$. Each voter of type $i$ will vote for party $A$ if and only if $\Delta_{i}(x_{A},x_{B})\geq\epsilon$, which holds with probability $\pi_{i}(x_{A},x_{B})=F(\Delta_{i}(x_{A},x_{B}))$. As standard in probabilistic voting models, we let the support of $F$ be wide, so that cases where $\pi_{i}\in\{0,1\}$ can be ignored in the equilibrium analysis. [^8] Formally, for $x_{\lor}=(\max_{i\in I}{x^{k}_{i}})_{k\leq K}$ and $x_{\land}=(\min_{i\in I}{x^{k}_{i}})_{k\leq K}$, we assume that both $v(x_{\land},x_{\lor})-v(x_{\lor},x_{\lor})$ and $v(x_{\lor},x_{\lor})-v(x_{\land},x_{\lor})$ are in the support of $F$.
@@
Part (i) requires that rents increase in the number of votes. [^9] Proposition 2.1 would generalize if $R$ was weakly increasing on $[0,\frac{1}{2}]$ and not constant on $[0,\frac{1}{2})$. In that case, party divergence would obtain for some distributions of voter types. We do not assume continuity: this allows, for instance, for jumps in the office rents as a party’s vote share meets institutional thresholds that grant agenda-setting power, the power to block legislation or entitlement to electoral reimbursements. In particular, the general setup allows for a party to gain a sizable share of power (and political rents) once it obtains the absolute majority of votes. An alternative interpretation of this setup (proposed, e.g., by Glaeser et al.,, 2005), is that voters are making a campaign contribution decision rather than a voting decision, so that parties’ rents naturally grow with their supporters. [^10] Under this interpretation, the size $s_{i}$ of voter group $i$ would already account for the heterogeneity in voters’ budget devoted to political contributions. Part (ii) imposes the restriction that aggregate rents are constant. Our second main assumption is that parties’ utility from rents is increasing and concave. Formally, let the payoff function of each party be $U:\mathbb{R}\to\mathbb{R}$. The following holds.
@@
Let $\nu:[0,1]\to\mathbb{R}$, $\nu=U\circ R$, be the function mapping a party’s vote share to its payoff, and without any loss of generality, let payoffs be normalized so that $\nu(1)-\nu(0)=1$. The following holds when assumptions 1 and 2 are combined.
@@ 2.1 Party Divergence
If a pure strategy equilibrium exists, candidates offer different policy platforms.
@@ 3 Unidimensional Conflict
We start by analyzing the case where $K=1$, so that parties compete on a single policy dimension. [^13] In this case, dimension indexes $k$ are redundant, and omitted. Let $\epsilon\sim U(-\phi,\phi$), so that the shock is uniformly distributed, and assume that $h(|x-x_{i}|)=-\alpha(x-x_{i})^{2}$, where $\alpha>0$ captures how important policies are for individual payoffs and, without loss of generality, we set $\alpha=1$. With one policy dimension, we can think of voter heterogeneous preferences as capturing the classical economy-based conflict of interest between right-wing and left-wing voters, as in the following simple example.
@@
For the next propositions, assume that voters are ordered so that $x_{1}\leq x_{2}\leq...\leq x_{N}$, and let $x_{m}$ be the median voter type. If the median voter does not exist, define $x_{m}$ to be the average of the two “most median” voter types. We show that the game admits a pure strategy equilibrium, and there are exactly two platforms that can be offered in such an equilibrium. One policy lies to the left of the risk-neutral equilibrium policy $x^{RN}$, where
$x^{RN}=\sum_{i=1}^{N}\frac{R(s_{i}+\sum_{j<i}s_{j})-R(\sum_{j<i}s_{j})}{\bar{R}}x_{i},$
with $s_{0}=0$. The other lies to the right of such policy.
Proposition 3.1 (Equilibrium)
The game has a pure strategy equilibrium. In particular:
  1. (i)
    There exist two platforms, $\underaccent{\bar}{x},\bar{x}\in\mathbb{R}$ with $x_{1}<\underaccent{\bar}{x}<x^{RN}<\bar{x}<x_{N}$, such that, in every pure strategy equilibrium, one candidate offers $\underaccent{\bar}{x}$ and the other offers $\bar{x}$. It has
@@
    where $\bar{w}_{i},\underaccent{\bar}{w}_{i}\in(0,1)$ for each $i\in I$, $\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=\sum_{i=1}^{N}\bar{w}_{i}=1$.
  2. (ii)
    For each $i\in I$, the equilibrium weights $\bar{w}_{i}$ and $\underaccent{\bar}{w}_{i}$ attributed to $x_{i}$ are proportional to their payoff increase from winning the votes of group $i$. In particular
    $\underaccent{\bar}{w}_{i}=\frac{\nu(\underaccent{\bar}{s}_{i})-\nu(\underaccent{\bar}{s}_{i-1})}{\nu(1)-\nu(0)},\quad\quad\quad\quad\bar{w}_{i}=\frac{\nu(\bar{s}_{i+1})-\nu(\bar{s}_{i})}{\nu(1)-\nu(0)},$
    where, for each $i\in I$, $\underaccent{\bar}{s}_{i}=\sum_{j\leq i}s_{j}$ and $\bar{s}_{i}=\sum_{j\geq i}s_{j}$.
  3. (iii)
    One candidate attributes more weight to the policy demands of those to the left of the median voter, the other candidate attributes more weight to the demands of those to the right of the median voter. In particular, for each $i\in I$,
    $x_{i}>(<)\text{ }x_{m}\implies\bar{w}_{i}>(<)\text{ }\underaccent{\bar}{w}^{i}$
    where $x_{m}$ is the median voter type.
  4. (iv)
    Finally, the equilibrium expected payoff of each party is proportional to the (squared) distance between the two equilibrium platforms $(\bar{x}-\underaccent{\bar}{x})^{2}$, up to a constant term.
@@ 3.1 Voter Alienation and Polarization
In other words, we say that candidate $A$ benefits from attracting voter group $i$ if a change in voter group $i$’s views that increase its support for candidate $A$ improves candidate $A$’s equilibrium payoff. Conversely, candidate A benefits from alienating voting group $i$ if an increase in group $i$’s support for party $A$ decreases party $A$’s equilibrium payoff.
At this point, one could be skeptical that a candidate ever benefits from alienating a group of voters, given that it means that the rest of the voter distribution remains unchanged and that group of voters will be less likely to support the party in the new equilibrium. Surprisingly, our next result suggests that each party always has the incentive to alienate at least one voter group: the most extremist supporters of the rival candidate.
@@ 4 Multidimensional Conflict
Let us now consider $K>1$ policies, so that platforms $x_{A}$ and $x_{B}$ are vectors of $\mathbb{R}^{K}$. While we still denote a voter distribution by $(\mathbf{x},\mathbf{s})$, note that now a type $x_{i}$ is also a vector of $\mathbb{R}^{K}$. The analysis is carried out under the same assumptions of section 3, with some additional restrictions. First, we require $\nu$ to be strictly concave. Strict concavity guarantees that party best responses never fall at points where payoffs are non-differentiable, which greatly simplifies the analysis. [^18] The assumption can be relaxed. In particular, we can allow for a jump in payoff when parties reach the majority of votes, provided that a median voter type exists. Second, for some of the main results of the section we restrict the attention to symmetric voter distribution, which we prove to admit a pure strategy Nash equilibrium. [^19] In the more general formulation of the game, the existence of a mixed strategy equilibrium is guaranteed by Glicksberg’s existence result, but the existence of a pure strategy equilibrium is not, because the utility function of a party is not quasi-concave in that party’s strategy.
@@ 4.1 Equilibrium Analysis
The equilibrium analysis presents some technical challenges. First, parties’ expected payoffs are not differentiable at platform profiles $(x_{A},x_{B})$ such that two voter groups, $i$ and $j$, have the same relative support for party $A$ (i.e., $\Delta_{i}(x_{A},x_{B})=\Delta_{j}(x_{A},x_{B})$). Second, a party’s expected payoff is not quasi-concave in that party platform – best responses are not continuous in the competitors’ platforms. Third, the game is not finite, in the sense that the strategy set is non-finite. These features complicate the search for pure strategy equilibria, as even their existence is not guaranteed by standard results. The first step of the equilibrium analysis is introducing the notion of local equilibrium of the election game.
@@ 4.1.1 Algorithmic Solutions
Notes. Panel (a) plots a non-symmetric bidimensional voter distribution and the associated local equilibria where party payoffs. Panel (b) plots a symmetric bidimensional voter distribution and the associated local equilibria. In both panels: (i) “x” markers represent voter types’ ideal policies, and each voter group is equally sized; (ii) two circles of the same color represent a pair of local equilibrium platforms; (iii) we assume that party’s payoff as a function of vote shares is $\nu(s)=\ln(\frac{1}{2}+s)$.
Figure 1 panel (a) reports a voter distribution over a bidimensional policy space, as well as the set of local equilibria, found applying the algorithm outlined in the previous paragraph. Lemma 4.1 tells us that among the existing local equilibria, the one identified by the two dark dots is associated to the highest party expected payoff since the platforms proposed by the two politicians differ the most. In panel (b) we report the voter types and local equilibria for a symmetric voter distribution. In this case, the local equilibrium identified by the two dark dots is also the Nash equilibrium where parties expect to benefit the most, as per the results of proposition 4.1.
@@ 4.2 Divide and Diverge
Figure 2 reports the direction in which voter opinions should change for parties’ equilibrium payoffs to increase the fastest, for the two local equilibria of figure 1, panel (a). The red arrows represent this direction for the best local equilibrium (identified by the dark circles), while the blue arrows identify the direction for the other local equilibrium (identified by the white circles). In both cases, the change makes supporters of party $A$ on average more different from supporters of party $B$.
@@ 5.1 Common Interest vs Conflict of Interest Issues
A recent stream of literature in political science and economics has stressed how politicians tend to focus on divisive, zero-sum, issues over common interest issues, even if the latter entail policies that could benefit a larger share of voters (e.g., Fiorina et al.,, 2011; Matsusaka and Kendall,, 2022; Bueno de Mesquita and Dziuda,, 2023). This phenomenon is potentially amplified by the fact that, on the one hand, voters seem to be prone to social and political “zero-sum thinking,” the perception that gains from one social group must necessarily come at the expense of another (e.g., Carvalho et al.,, 2023; Chinoy et al.,, 2023; Ali et al.,, 2024), and often lack the information to assess the “common value” nature of policies (Matsusaka and Kendall,, 2022), while politicians’ rhetoric often paints politics in a zero-sum fashion. Our simple model can provide a new perspective on why politicians might have an incentive to focus on more divisive issues, stress the zero-sum component of policy-making, and keep voters informed on zero-sum issues and uninformed about common interest policies.
@@ 6 Some Words on Welfare
so that, when $R_{m}\to\bar{R}$, $\bar{x}\to x^{o}$ and $\underaccent{\bar}{x}\to x^{o}$. This result is captured by the following remark.
@@ A.1 Proofs
We start from part (i). Fix a distribution $(\mathbf{x},\mathbf{s})$ and order $i,...,N$ so that $x_{1}\leq x_{2}\leq...\leq x_{N}$. Fix $p\in\{A,B\}$ and let $x_{-p}\in\mathbb{R}$. As a first step we show that, if $x_{p}$ is a best response to $x_{-p}$, then either $x_{p}>x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\bar{w}_{i}x_{i}$, for $\bar{w}_{i}=\frac{\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})}{\nu(1)-\nu(0)}$, or $x_{p}<x_{-p}$ and $x_{p}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}x_{i}$, for $\underaccent{\bar}{w}_{i}=\frac{\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})}{\nu(1)-\nu(0)}$.
By proposition 2.1, we know that for any platform $x\in\mathbb{R}$, $x$ is never the best response to $x$. Assume that $x_{p}$ is a best response to $x_{-p}$. Given that $x_{p}\neq x_{-p}$ there are two cases. The first is $x_{p}>x_{-p}$. Subject to the constraint $x_{p}>x$, and fixing $x_{-p}$,$V_{p}$ is strictly concave in $p$’s platform $x$. Hence, $x_{p}$ it must satisfy the first-order condition $\frac{\partial V_{p}}{\partial x_{p}}=0$. When $p=A$, the first order condition is
@@
Next, we show that for each $p\in\{A,B\}$ and each $x_{B}\in\mathbb{R}$, $A$ has at least one best response to any $x_{B}$. Fix $x_{B}\in\mathbb{R}$. First, note that any platform choice outside $[x_{1},x_{N}]$ is strictly dominated by either $x_{1}$ or $x_{N}$. We can then focus on looking for best responses in $[x_{1},x_{N}]$. Next, note that $V_{A}(\cdot,x_{B})$ is continuous on $[x_{1},x_{N}]$. Hence, by the extreme value theorem, for any $x_{B}$, $A$ has at least one best response to $x_{B}$. For the same reason, $B$ has at least one best response to any $x_{A}\in\mathbb{R}$. Given that the best response correspondence is never empty valued, to show that a pure strategy Nash equilibrium exists it suffices to show that $\underaccent{\bar}{x}<\bar{x}$.
First, note that since $\underaccent{\bar}{w}_{i},\bar{w}_{i}\in[0,1]$ for $i=1,...,N$, and $\sum_{i=1}^{N}\bar{w}_{i}=\sum_{i=1}^{N}\underaccent{\bar}{w}_{i}=1$, $(\bar{w}_{i},x_{i})_{i\in I}$ and $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$ are probability distributions. Second, $(\bar{w}_{i},x_{i})_{i\in I}$ strictly first order dominates $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$, as we now show. Since the probability distributions share support $\{x_{1},...,x_{N}\}$, it is sufficient to show that, for each $i=1,...,N$
@@
Finally, note that $\bar{x}$ is the expected value of $(\bar{w}_{i},x_{i})_{i\in I}$ and $\underaccent{\bar}{x}$ is the expected value of $(\underaccent{\bar}{w}_{i},x_{i})_{i\in I}$. Strict first order stochastic dominance implies $\bar{x}>\underaccent{\bar}{x}$. This concludes the proof that the pure strategy equilibrium exists, and that only $\bar{x}$ and $\underaccent{\bar}{x}$ will be played in such equilibrium.
To prove part (ii), note that for each $i\in I$, it has $\bar{w}_{i}=\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})$ and $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$. The former weight represents the payoff gain of the party offering $\bar{x}$ when she wins type $i$. Voter group $i$ will vote for the party offering $\bar{x}$ only if all $j>i$ – that is, a share $\bar{s}_{i+1}$ of the distribution, also vote for that party. If voter group $i$ also votes for the party offering $\bar{x}$, the vote share of that party increases from $\bar{s}_{i+1}$ to $\bar{s}_{i}=\bar{s}_{i+1}+s_{i}$, with an increase in payoff equal to $\nu(\bar{s}_{i})-\nu(\bar{s}_{i+1})=\bar{w}_{i}$. Using the same argument, it is easy to show that $\underaccent{\bar}{w}_{i}=\nu(1-\bar{s}_{i+1})-\nu(1-\bar{s}_{i})$ is the payoff gain of the party offering $\underaccent{\bar}{x}$ from winning voter group $i$.
Part (iii) follows immediately. Note that, by remark 1, it has